LASER-INDUCED FLUORESCENCE SPECTROSCOPY ON ROTATIONAL DISTRIBUTION OF HfF PHO-TOIONS

\underline{MATT GRAU}, HUANQIAN LOH, TYLER YAHN, RUSSELL STUTZ, JILA, NIST and University of Colorado, and Department of Physics, University of Colorado, Boulder, Colorado 80309-0440; ROBERT W. FIELD, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, Massachusetts 02139; and ERIC A. CORNELL, JILA, NIST and University of Colorado, and Department of Physics, University of Colorado, Boulder, Colorado 80309-0440.

As a step towards measuring the electron electric dipole moment^{a}, we produce a sample of HfF^{+} using a two-color excitation. We promote HfF from X^{2}Δ_{3/2} to an isotope and parity-selective intermediate state, and then to one of many highly perturbed Rydberg states from which it autoionizes to the vibrational ground state of HfF^{+}. We measure the population of the rotational states of HfF^{+} using laser-induced fluorescence and find that only a small number of states are populated, with most of the population in J < 4. Additionally, we see a strong propensity for autoionization to preserve the parity of the molecule, with one parity populating even J levels and the other populating odd J. Using polarized light to prepare the Rydberg molecules in various orientations, and then probing the ion with LIF, we see that a polarization of m_{J} sublevels also survives autoionization.

aA. Leanhardt et al, arXiv:atom-ph/1008.2997v2 (2010).